OpenStudy (4itc):

(4/√2) + (2/√3) = ? a) (4√3 + 2√2) / (√5) b) (4√3 + 2√2) / (√6) c) (6) / (√2 + √3) d) (6) / (√5) e) (8) / (√6)

OpenStudy (4itc):

i chose d. but i think its wrong

OpenStudy (plainntall):

\[\frac{ 4 }{ \sqrt{2} }+\frac{ 2 }{ \sqrt{3} }\]

OpenStudy (4itc):

i added them straight across... not sure if the same denominator still works if they are roots

OpenStudy (plainntall):

Get a common denominator \[\frac{ 4\sqrt{3} }{ \sqrt{2}\sqrt{3}}+\frac{ 2\sqrt{2} }{ \sqrt{2}\sqrt{3}}\]

OpenStudy (welshfella):

the LCD of sqrt3 and sqrt 2 is sqrt 6 to get the denominator multiply each term by sqrt 6

OpenStudy (plainntall):

Can you see where this is going yet?

OpenStudy (4itc):

well the denominators equal √5 and √5... i dont think you add them. Yea! no adding! the numerators dont change. sooo B.

OpenStudy (plainntall):

\[\frac{ 4\sqrt{3}+2\sqrt{2} }{ \sqrt{6} }\] based on \[\sqrt{a}\sqrt{b}=\sqrt{ab}\] I don't think they want you to crunch numbers. I think they want you to use your understanding of how products of square roots relate.

OpenStudy (4itc):

ohhhh times not add.

OpenStudy (4itc):

thanks you guys!

OpenStudy (plainntall):

:)